OpenStudy (anonymous):

Use the compound interest formula to determine the final value of the following amount. $1300 at 11.9% compounded monthly for 2.5 years. What is the value of the amount?

OpenStudy (anonymous):

I'm thinking something like A=1300(1.119/12)^12*2.5 ???

OpenStudy (anonymous):

\[I = P(1 + \frac rn)^{nt}\] Where: P = Principal r = rate in decimal form n = number of times compounded in a year t = number of years

OpenStudy (anonymous):

What goes in the parenthesis? (1+12/.119x12)

OpenStudy (anonymous):

One thing is off. Your term inside the parentheses. Inside the parentheses is 1 + 0.119/12 which is 1 + 0.00991666...

OpenStudy (anonymous):

Can you take it from here?

OpenStudy (anonymous):

So... A=13000(1.009916667)^30 So... 17478.72??

OpenStudy (anonymous):

I got 1747.8716

OpenStudy (anonymous):

I typed 13000

OpenStudy (anonymous):

There's the problem :)

OpenStudy (anonymous):

Ugh...these are terrible. Thank you very much

OpenStudy (anonymous):

You're welcome :)